# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 2.1.1. PEL datum
  • $B$ is a finite dimensional semisimple algebra over $\mathbb{Q}$ with a positive involution ***. We further assume that $B_{\mathbb{Q}_{p}}$ is isomorphic to a product of matrix algebras over finite extensions of $\mathbb{Q}_{p}$ (note that we allow the extensions ramified over $\mathbb{Q}_{p}$).
@@
  • $\left\langle\cdot,\cdot\right\rangle:V\times V\to\mathbb{Q}$ is a symplectic form on $V$ such that $\left\langle bv,w\right\rangle=\left\langle v,b^{*}w\right\rangle$ for all $v,w\in V$ and $b\in B$.
  • $\operatorname{\mathcal{O}}_{B}$ is a ***-invariant $\mathbb{Z}_{(p)}$-order of $B$ such that $\operatorname{\mathcal{O}}_{B}\otimes\mathbb{Z}_{p}$ is a maximal $\mathbb{Z}_{p}$-order of $B_{\mathbb{Q}_{p}}$.
[…]
@@
  • $h:\mathbb{S}=\operatorname{Res}_{\mathbb{C}|\mathbb{R}}(\mathbb{G}_{m})\to G_{\mathbb{R}}$ is a group homomorphism which defines a Hodge structure of type $\{(-1,0),(0,-1)\}$ on $V_{\mathbb{R}}$ and the form $\left\langle\cdot,h(\sqrt{-1})\cdot\right\rangle$ is symmetric and positive definite.
@@
We make some explicit description of this datum, which will be fixed in the rest of this paper. By assumption, there is a decomposition (as product of ***-invariant simple factors)
@@
where for each $1\leq i\leq r$, $R_{i}$ is a product of finite extensions (may be ramified) of $\mathbb{Q}_{p}$. Let $I:=\{1,\dots,r\}$ be the index set. For each $i\in I$, let $F_{i}$ be the set of ***-invariant elements in $R_{i}$, which is a field over $\operatorname{\mathbb{Q}}_{p}$. Since we require the group $G$ to be connected, we can decompose the index set $I$ into four types: (C), (AL), (AU), (AR) (see also [4] subsection 2.2, but note that our $I$ is the quotient of their $\{1,\dots,r\}$ by the induced action of $\ast$). For each $i\in I$, its type is defined by
@@
    if *** induces the identity on $R_{i}$, so we have $R_{i}={F_{i}}$.
@@
    if $R_{i}\simeq{F_{i}}\times{F_{i}}$ and *** exchanges the two factors.
@@
    if *** is an automorphism of order 2 on $R_{i}$, and $R_{i}/F_{i}$ an unramified quadratic extension.
@@
    if *** is an automorphism of order 2 on $R_{i}$, and $R_{i}/F_{i}$ a ramified quadratic extension.
@@
For each $i\in I$, the existence of ***-symplectic form implies that $V_{i}$ is a vector space over $F_{i}$ of even dimension, so we write
@@
where each $\Lambda_{i}$ is an $\operatorname{\mathcal{O}}_{R_{i}}$-lattice in $V_{i}$, and there is an induced pairing $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i}\times\Lambda_{i}\rightarrow\mathbb{Z}_{p}$, which factors through $\operatorname{\mathcal{O}}_{F_{i}}$.
@@ 2.1.2. Rational group structure
[…]
@@
is the subgroup such that the local similitude factors $c_{i}$ are the same and defined over $\mathbb{Q}_{p}$. From now on, we fix an isomorphism of fields ¯$\mathbb{C}\simeq\overline{\mathbb{Q}}_{p}$. Then we view $\mu$ as a cocharacter of $G$ over ¯$\overline{\mathbb{Q}}_{p}$. Let $F$ be a fixed sufficiently large finite Galois extension of $\mathbb{Q}_{p}$ containing all embeddings of $F_{i}$ to ¯$\overline{\mathbb{Q}}_{p}$ such that $G_{F}$ is split. Then the cocharacter $\mu$ is defined over $F$ and we can write
@@
[…]
@@
Since the only weights of $\mu:\mathbb{G}_{m,F}\to G_{F}$ are 00 and $1$, we get an induced decomposition
@@
[…]
@@
Similarly for $W^{\prime}$. For each $i,j,l$, we have $V_{i,j}^{l}=W_{i,j}^{l}\oplus(W^{\prime})_{i,j}^{l}$. Let $d_{i,j}^{l}=\dim_{F}W_{i,j}^{l}$. If we write $d_{i,j}^{l}=d_{\sigma_{i,j}^{l}}$ to make explicit the embedding $\sigma_{i,j}^{l}$, then there is an induced ***-action on $\{\sigma_{i,j}^{l}\}_{i,j,l}$ and we set $(d_{i,j}^{l})^{*}:=d_{(\sigma_{i,j}^{l})^{*}}$. The natural isomorphism $W\simeq(W^{\prime})^{\vee}$ gives the identity
@@
In the rest of this paper, the index $(i,j,l)$ will always be the tuple of integers runs through $i\in I,1\leq j\leq f_{i}$ and $0\leq l\leq e_{i}$.
@@ 2.1.3. Integral group structure
Now we discuss the integral group theoretic data which will be needed later. Consider the parahoric group scheme $\mathcal{G}$ over $\mathbb{Z}_{p}$ defined by the integral PEL datum $(\operatorname{\mathcal{O}}_{B},\ast,\Lambda,\langle\cdot,\cdot\rangle)$. Recall the decomposition $\Lambda=\bigoplus_{i=1}^{r}\Lambda_{i}^{m_{i}}$, where each $\Lambda_{i}$ is an $\operatorname{\mathcal{O}}_{R_{i}}$-lattice in $V_{i}$ together with an induced form $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i}\times\Lambda_{i}\rightarrow\operatorname{\mathcal{O}}_{F_{i}}$. Let $\operatorname{\mathcal{G}}_{i}/\operatorname{\mathbb{Z}}_{p}$ be the parahoric subgorup of $G_{i}$ associated to $(\Lambda_{i},\langle\cdot,\cdot\rangle_{i})$, then
@@
which admits an induced $\kappa_{i}[\varepsilon_{i}]$-action and a pairing $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i,0}\times\Lambda_{i,0}\rightarrow\kappa_{i}$. Let $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$ be the maximal reductive quotient of $\operatorname{\mathcal{G}}_{0}$. We have similarly an inclusion
@@ 2.1.4. The group $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$
For each $i\in I$, we define the group $\operatorname{\mathcal{G}}_{i,0}^{\mathrm{spl}}/\operatorname{\mathbb{F}}_{p}$ as similitude automorphism group associated to $(\Lambda_{i,0}^{\operatorname{rdt}})^{e_{i}}$ with its natural pairing. Then we have an isomorphism
@@
The pairing on $\Lambda_{i}$ induces a natural pairing on the $\operatorname{\mathcal{O}}_{F}$-lattice $\Lambda_{i,j}^{l}$, and the similitude automorphism group is isomorphic to $G_{i,j}^{l}$ over $\operatorname{\mathcal{O}}_{F}$. In contrast to the rational case, in general $\operatorname{\mathcal{G}}_{i,\operatorname{\mathcal{O}}_{F}}$ does not split as product of $G_{i,j}^{l}$ due to ramification. So we define the split $\operatorname{\mathcal{O}}_{F}$-lattice as
@@
by requiring having the same $\mathbb{Z}_{p}$-local similitude factors. Over $F$ we have the following isomorphism
@@
In other words, $\mathcal{G}^{\mathrm{spl}}_{0}$ is an $\mathbb{F}_{p}$ model of the special fiber of $\mathcal{G}^{\mathrm{spl}}$. Moreover, one can analysis the decomposition of $\Lambda_{i,0}^{\mathrm{spl}}$ induced by $\mu_{i}$ exactly as in the characteristic zero case (see the end of in 2.1.2).
@@ 2.2. Smooth Pappas-Rapoport splitting models
Recall that we have $\mu:\mathbb{G}_{m\overline{\mathbb{Q}}_{p}}\rightarrow G_{\overline{\mathbb{Q}}_{p}}$ the Hodge cocharacter of $G$ over $\overline{\mathbb{F}}_{p}$. Let $E$ be the field of definition of the conjugacy class attached to $\mu$. Then by the assumption on $F$, we have $E\subset F$.
Let $K^{p}$ be a sufficiently small open compact subgroup of $G(\mathbb{A}_{f}^{p})$, which will be fixed in the rest of this section. By works of Kottwitz [28] and Rapoport-Zink [53], there is a scheme $\mathscr{A}^{\operatorname{naive}}=\mathscr{A}^{\operatorname{naive}}_{K^{p}}$ over $\operatorname{\mathcal{O}}_{E}$ representing the following moduli problem (see [53] chapter 6 for more details): for any $\operatorname{\mathcal{O}}_{E}$-scheme $S$, $\mathscr{A}^{\operatorname{naive}}(S)$ classifies the isogeny class of tuples $(A,\lambda,\iota,\alpha)$, where
@@
where $\omega_{A/S}$ is the sheaf of invariant differentials. For simplicity, write $\omega=\omega_{A/S}$, and $\operatorname{\mathcal{H}}=H^{1}_{\operatorname{dR}}(A/S)$. The action of $\operatorname{\mathcal{O}}_{B}\otimes\mathbb{Z}_{p}$ on the $\operatorname{\mathcal{O}}_{S}$-modules $\operatorname{\mathcal{H}}$ and $\omega$ induces decompositions
@@
$0\to\omega_{i,j}\to\mathcal{H}_{i,j}\to\operatorname{Lie}_{A^{\vee}/S,i,j}\to 0$
@@
$Q_{i,j}^{\leq l}(T)=\prod_{k=1}^{l}(T-\pi_{i,j}^{k})\quad\text{ and }\quad Q_{i,j}^{>l}=\prod_{k=l+1}^{e_{i}}(T-\pi_{i,j}^{k}).$
@@
Let $S$ be an $\operatorname{\mathcal{O}}_{F}$-scheme and $\underline{A}=(A,\iota,\lambda,\alpha)\in\mathscr{A}^{\operatorname{naive}}(S)$ with the associated $\mathcal{H}$ and $\omega$ together with the induced additional structure. A splitting structure with respect to $\underline{A}$ is given by the datum ¯$\underline{\mathscr{F}_{\bullet}}=(\mathscr{F}_{i,j}^{l})$: for each $i\in I,1\leq j\leq f_{i}$, there is a filtration of $\mathcal{O}_{S}$-modules with $\operatorname{\mathcal{O}}_{F_{i}}$-action
@@
    Write $[\pi_{i}]$ for the action of $\pi_{i}\in\operatorname{\mathcal{O}}_{F_{i}}$. Then for all $1\leq l\leq e_{i}$, we have
@@
    $([\pi_{i}]-\pi_{i,j}^{e_{i}})\cdots([\pi_{i}]-\pi_{i,j}^{e_{i}+l-1})\mathscr{F}_{i,j}^{e_{i}+l}\subset\mathscr{F}_{i,j}^{e_{i}-l}.$
More explicitly, we have the following description of splitting structures according to the type of $i$ (see the notations in 2.1.2; see also [4] subsection 2.3):
@@
    where for each $1\leq l\leq e_{i}$, $\operatorname{\mathscr{F}}_{i,j}^{l}$ is a locally a direct factor of rank $d_{i}l$ such that
    $\operatorname{\mathscr{F}}_{i,j}^{l,\perp}=(\pi_{i})^{e_{i}-l}(\operatorname{\mathscr{F}}_{i,j}^{l})$
    taken in $\operatorname{\mathcal{H}}_{i,j}$.
  • (AL): We have $R_{i}=F_{i}\times F_{i}$, and $\operatorname{\mathcal{H}}_{i}=\operatorname{\mathcal{H}}_{i}^{\prime}\times\operatorname{\mathcal{H}}_{i}^{{}^{\prime}\vee}$. For each $j$, the splitting structure is reduced to a filtration
@@
    where each $\operatorname{\mathscr{F}}_{i,j}^{l}$ is locally a direct factor of rank $a_{i,j}^{1}+\cdots+a_{i,j}^{l}$.
@@
    $\omega_{i,j}=\omega_{i,j}^{\prime}\times(\omega_{i,j}^{\prime})^{\vee}$
    given by the unramified $\operatorname{\mathcal{O}}_{R_{i}}$ action, and the splitting structure is reduced to a filtration
@@
    where each $\operatorname{\mathscr{F}}_{i,j}^{l}$ is locally a direct factor of rank $a_{i,j}^{1}+\cdots+a_{i,j}^{l}$.
@@
  • ¯$\underline{\mathscr{F}_{\bullet}}=(\mathscr{F}_{i,j}^{l})$ is a splitting structure of $\underline{A}$.
  • The isomorphism is given by $\mathbb{Z}_{(p)}^{\times}$-isogenies between abelian schemes with $\operatorname{\mathcal{O}}_{B}\otimes_{\mathbb{Z}}\mathbb{Z}_{(p)}$-structures and splitting structure.
@@
By Propositions 5.2 and 9.2 of [47], when $i$ is of type (AL) or type (C), such isomorphism exists locally. Moreover, when $i$ is of type (AU), the same proof equally applies.
@@
which is $\prod_{i,j,l}G_{i,j}^{l}$-equivariant. The morphism $\pi$ is the natural forgetful functor, which is a $\prod_{i,j,l}G_{i,j}^{l}$-torsor.
@@
is a resolution of singularities. Note that in this case, the associated parahoric subgroup $K_{p}=\operatorname{\mathcal{G}}(\mathbb{Z}_{p})$ is very special in the sense of [68] Definition 6.1. The purpose of this paper is to show that in our ramified setting, the smooth splitting model $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ admits many nice properties as in the unramified setting. Recall for any $(V,\eta)\in\mathrm{Rep}_{F}P_{\mu}$, we have the associated automorphic vector bundle $\mathcal{V}$ over
@@
cf. [39]. Since $G$ is split over $F$, we have a canonical reductive model $G$ together with the parabolic $P_{\mu}$ over $\operatorname{\mathcal{O}}_{F}$, and $(V,\eta)$ naturally extends to a representation of $P_{\mu}$ over $\operatorname{\mathcal{O}}_{F}$. In general, it is hard to extend $\mathcal{V}$ to an vector bundle on $\operatorname{\mathscr{A}}_{\operatorname{\mathcal{O}}_{F}}$. Nevertheless, we have

### v3

@@ 2.1.1. PEL datum
  • $B$ is a finite dimensional semisimple algebra over $\mathbb{Q}$ with a positive involution *. We further assume that $B_{\mathbb{Q}_{p}}$ is isomorphic to a product of matrix algebras over finite extensions of $\mathbb{Q}_{p}$ (note that we allow the extensions to be ramified over $\mathbb{Q}_{p}$).
@@
  • $\left\langle\cdot,\cdot\right\rangle:V\times V\to\mathbb{Q}$ is a non-degenerate symplectic form on $V$ such that $\left\langle bv,w\right\rangle=\left\langle v,b^{*}w\right\rangle$ for all $v,w\in V$ and $b\in B$.
  • $\operatorname{\mathcal{O}}_{B}$ is a *-invariant $\mathbb{Z}_{(p)}$-order of $B$ such that $\operatorname{\mathcal{O}}_{B}\otimes\mathbb{Z}_{p}$ is a maximal $\mathbb{Z}_{p}$-order of $B_{\mathbb{Q}_{p}}$.
  • $\Lambda$ is a $\operatorname{\mathcal{O}}_{B}\otimes\mathbb{Z}_{p}$-lattice in $V_{\mathbb{Q}_{p}}$, such that $\left\langle\cdot,\cdot\right\rangle$ induces a (not necessary perfect) pairing $\Lambda\times\Lambda\to\mathbb{Z}_{p}$.
@@
  • $h:\mathbb{S}=\operatorname{Res}_{\mathbb{C}|\mathbb{R}}(\mathbb{G}_{m})\to G_{\mathbb{R}}$ is a group homomorphism which defines a Hodge structure of type $\{(-1,0),(0,-1)\}$ on $V_{\mathbb{R}}$ and the form $\left\langle\cdot,h(\sqrt{-1})\cdot\right\rangle$ is symmetric and positive definite. (Note that $(\Lambda,\left\langle\cdot,\cdot\right\rangle,h)$ is a PEL-type $\operatorname{\mathcal{O}}_{B}$-lattice in the sense of [33] Definition 1.2.1.3.)
@@
We make some explicit description of this datum, which will be fixed in the rest of this paper. By assumption, there is a decomposition (as product of *-invariant simple factors)
@@
where for each $1\leq i\leq r$, $R_{i}$ is a product of finite extensions (maybe ramified) of $\mathbb{Q}_{p}$. Let $I:=\{1,\dots,r\}$ be the index set. For each $i\in I$, let $F_{i}$ be the set of *-invariant elements in $R_{i}$, which is a field over $\operatorname{\mathbb{Q}}_{p}$. Since we require the group $G$ to be connected, we can decompose the index set $I$ into four types: (C), (AL), (AU), (AR) (see also [6] subsection 2.2, but note that our $I$ is the quotient of their $\{1,\dots,r\}$ by the induced action of $\ast$). For each $i\in I$, its type is defined by
@@
    if * induces the identity on $R_{i}$, so we have $R_{i}={F_{i}}$.
@@
    if $R_{i}\simeq{F_{i}}\times{F_{i}}$ and * exchanges the two factors.
@@
    if * is an automorphism of order 2 on $R_{i}$, and $R_{i}|F_{i}$ an unramified quadratic extension.
@@
    if * is an automorphism of order 2 on $R_{i}$, and $R_{i}|F_{i}$ a ramified quadratic extension.
@@
For each $i\in I$, the existence of *-symplectic form implies that $V_{i}$ is a vector space over $F_{i}$ of even dimension, so we write
@@
where each $\Lambda_{i}$ is an $\operatorname{\mathcal{O}}_{R_{i}}$-lattice in $V_{i}$, and there is an induced pairing $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i}\times\Lambda_{i}\rightarrow\operatorname{\mathcal{O}}_{R_{i}}$ if $i$ is of type (C) or (AL), and $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i}\times\Lambda_{i}\rightarrow\operatorname{\mathcal{O}}_{F_{i}}$ if $i$ is of type (AU).
@@ 2.1.2. Rational group structure
The restriction of the pairing $\langle\cdot,\cdot\rangle$ on $V_{i}$ defines an algebraic group $G_{i}=\operatorname{Res}_{F_{i}|\mathbb{Q}_{p}}H_{i}$ over $\operatorname{\mathbb{Q}}_{p}$: for any $F_{i}$-algebra $R$, we have
@@
is the subgroup such that the local similitude factors $c_{i}$ are the same and defined over $\mathbb{Q}_{p}$. From now on, we fix an isomorphism of fields $\mathbb{C}\simeq\overline{\mathbb{Q}}_{p}$. Then we view $\mu$ as a cocharacter of $G$ over $\overline{\mathbb{Q}}_{p}$. Let $F$ be a fixed sufficiently large finite Galois extension of $\mathbb{Q}_{p}$ containing all embeddings of $F_{i}$ to $\overline{\mathbb{Q}}_{p}$ such that $G_{F}$ is split. Then the cocharacter $\mu$ is defined over $F$ and we can write
@@
and the pairing $\langle\cdot,\cdot\rangle$ on $V_{i}$ induces pairings $\langle\cdot,\cdot\rangle:V_{i,j}^{l}\times V_{i,j}^{l}\rightarrow F$ (which in fact are defined over $F_{i}$). For each $i\in I$, as $G_{i}=\operatorname{Res}_{F_{i}|\operatorname{\mathbb{Q}}_{p}}H_{i}$, we have
@@
Since the only weights of $\mu:\mathbb{G}_{m,F}\to G_{F}$ are 0 and $1$, we get an induced decomposition
@@
where $z\in F^{\times}$ acts on $W$ (resp. $W^{\prime}$) by 1 (resp. $z$). This implies that the pairing $\langle\cdot,\cdot\rangle$ induces an isomorphism
@@
We have similar decompositions for $W^{\prime}$. For each $i,j,l$, we have $V_{i,j}^{l}=W_{i,j}^{l}\oplus(W^{\prime})_{i,j}^{l}$. Let $d_{i,j}^{l}=\dim_{F}W_{i,j}^{l}$. If we write $d_{i,j}^{l}=d_{\sigma_{i,j}^{l}}$ to make explicit the embedding $\sigma_{i,j}^{l}$, then there is an induced *-action on $\{\sigma_{i,j}^{l}\}_{i,j,l}$ and we set $(d_{i,j}^{l})^{*}:=d_{(\sigma_{i,j}^{l})^{*}}$. The natural isomorphism $W\simeq(W^{\prime})^{\vee}$ gives the identity
@@
In the rest of the paper, the index $(i,j,l)$ will always be a tuple of integers running through $i\in I,1\leq j\leq f_{i}$ and $0\leq l\leq e_{i}$.
@@ 2.1.3. Integral group structure
Now we discuss the integral group theoretic data which will be needed later. Consider the parahoric group scheme $\mathcal{G}$ over $\mathbb{Z}_{p}$ defined by the integral PEL datum $(\operatorname{\mathcal{O}}_{B},\ast,\Lambda,\langle\cdot,\cdot\rangle)$. Recall the decomposition $\Lambda=\bigoplus_{i=1}^{r}\Lambda_{i}^{m_{i}}$, where each $\Lambda_{i}$ is an $\operatorname{\mathcal{O}}_{R_{i}}$-lattice in $V_{i}$ together with an induced form $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i}\times\Lambda_{i}\rightarrow\operatorname{\mathcal{O}}_{R_{i}}$ (ii of type (C) or (AL)) or $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i}\times\Lambda_{i}\rightarrow\operatorname{\mathcal{O}}_{F_{i}}$ (ii of type (AU)). Let $\operatorname{\mathcal{G}}_{i}/\operatorname{\mathbb{Z}}_{p}$ be the parahoric subgroup of $G_{i}$ associated to $(\Lambda_{i},\langle\cdot,\cdot\rangle_{i})$, then
@@
which admits an induced $\kappa_{i}[\varepsilon_{i}]$-action and a pairing $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i,0}\times\Lambda_{i,0}\rightarrow\kappa_{i}$ (for $i$ of type (C) or (AU)) or $\langle\cdot,\cdot\rangle_{i}:\Lambda_{i,0}\times\Lambda_{i,0}\rightarrow\kappa_{i}\times\kappa_{i}$ (for $i$ of type (AL)). Let $\operatorname{\mathcal{G}}_{0}^{\operatorname{rdt}}$ be the maximal reductive quotient of $\operatorname{\mathcal{G}}_{0}$. We have similarly an inclusion
@@ 2.1.4. The group $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$
For each $i\in I$, we define the group $\operatorname{\mathcal{G}}_{i,0}^{\mathrm{spl}}/\operatorname{\mathbb{F}}_{p}$ as the similitude automorphism group associated to $(\Lambda_{i,0}^{\operatorname{rdt}})^{e_{i}}$ with its natural pairing. Then we have an isomorphism
@@
The pairing on $\Lambda_{i}$ induces a natural pairing on the $\operatorname{\mathcal{O}}_{F}$-lattice $\Lambda_{i,j}^{l}$, and the similitude automorphism group is isomorphic to $G_{i,j}^{l}$ over $\operatorname{\mathcal{O}}_{F}$. In contrast to the rational case, in general $\operatorname{\mathcal{G}}_{i,\operatorname{\mathcal{O}}_{F}}$ does not split as a product of $G_{i,j}^{l}$ due to ramification. So we define the split $\operatorname{\mathcal{O}}_{F}$-lattice as
@@
by requiring having the same $\mathbb{Z}_{p}^{\times}$-local similitude factors. Over $F$ we have the following isomorphism
@@
In other words, $\mathcal{G}^{\mathrm{spl}}_{0}$ is an $\mathbb{F}_{p}$ model of the special fiber of $\mathcal{G}^{\mathrm{spl}}$. Moreover, one can analyze the decomposition of $\Lambda_{i,0}^{\mathrm{spl}}$ induced by $\mu_{i}$ exactly as in the characteristic zero case (see the end of in 2.1.2).
@@ 2.2. Smooth Pappas-Rapoport splitting models
Recall that we have $\mu:\mathbb{G}_{m,\overline{\mathbb{Q}}_{p}}\rightarrow G_{\overline{\mathbb{Q}}_{p}}$ the Hodge cocharacter of $G$ over $\overline{\mathbb{Q}}_{p}$. Let $E$ be the field of definition of the conjugacy class attached to $\mu$. Then by the assumption on $F$, we have $E\subset F$.
Let $K^{p}$ be a sufficiently small open compact subgroup of $G(\mathbb{A}_{f}^{p})$, which will be fixed in the rest of this section. By works of Kottwitz [30] and Rapoport-Zink [55], there is a scheme $\mathscr{A}^{\operatorname{naive}}=\mathscr{A}^{\operatorname{naive}}_{K^{p}}$ over $\operatorname{\mathcal{O}}_{E}$ representing the following moduli problem (see [55] chapter 6 for more details): for any $\operatorname{\mathcal{O}}_{E}$-scheme $S$, $\mathscr{A}^{\operatorname{naive}}(S)$ classifies the isogeny classes of tuples $(A,\lambda,\iota,\alpha)$, where
@@
[…]
@@
$0\to\omega_{i,j}\to\mathcal{H}_{i,j}\to\operatorname{Lie}_{A^{\vee}/S,i,j}\to 0.$
@@
$Q_{i,j}^{\leq l}(T)=\prod_{k=1}^{l}(T-\pi_{i,j}^{k})\quad\text{ and }\quad Q_{i,j}^{>l}(T)=\prod_{k=l+1}^{e_{i}}(T-\pi_{i,j}^{k}).$
@@
Let $S$ be an $\operatorname{\mathcal{O}}_{F}$-scheme and $\underline{A}=(A,\iota,\lambda,\alpha)\in\mathscr{A}^{\operatorname{naive}}(S)$ with the associated $\mathcal{H}$ and $\omega$ together with the induced additional structure. A splitting structure with respect to $\underline{A}$ is given by the datum $\underline{\mathscr{F}_{\bullet}}=(\mathscr{F}_{i,j}^{l})$: for each $i\in I,1\leq j\leq f_{i}$, there is a filtration of $\mathcal{O}_{S}$-modules with $\operatorname{\mathcal{O}}_{F_{i}}$-action
@@
    Write $\pi_{i}]$ for the action of $\pi_{i}\in\operatorname{\mathcal{O}}_{F_{i}}$. Then for all $1\leq l\leq e_{i}$, we have
@@
    $\mathscr{F}_{i,j}^{e_{i}+l}=\mathscr{F}_{i,j}^{e_{i}-l,\perp}\quad\textrm{and}\quad([\pi_{i}]-\pi_{i,j}^{e_{i}})\cdots([\pi_{i}]-\pi_{i,j}^{e_{i}+l-1})\mathscr{F}_{i,j}^{e_{i}+l}\subset\mathscr{F}_{i,j}^{e_{i}-l}.$
    Here $\mathscr{F}_{i,j}^{e_{i}-l,\perp}$ is the orthogonal complement of $\mathscr{F}_{i,j}^{e_{i}-l}$ in $\operatorname{\mathcal{H}}_{i,j}$ under the pairing on $\operatorname{\mathcal{H}}_{i,j}$ induced by the polarization $\lambda$.
More explicitly, we have the following description of splitting structures according to the type of $i$ (see the notations in 2.1.2; see also [6] subsections 2.3 and 2.4):
@@
    where for each $1\leq l\leq e_{i}$, $\operatorname{\mathscr{F}}_{i,j}^{l}$ is a locally a direct factor of rank $d_{i}l$, and the filtration satisfies the relations as in the above conditions (2) and (4).
  • (AL): We have $R_{i}=F_{i}\times F_{i}$, and $\operatorname{\mathcal{H}}_{i}=\operatorname{\mathcal{H}}_{i}^{\prime}\oplus\operatorname{\mathcal{H}}_{i}^{{}^{\prime}\vee}$. For each $j$, the splitting structure is reduced to a filtration
@@
    where each $\operatorname{\mathscr{F}}_{i,j}^{l}$ is locally a direct factor of rank $a_{i,j}^{1}+\cdots+a_{i,j}^{l}$, and the filtration satisfies the above condition (2).
@@
    $\omega_{i,j}=\omega_{i,j}^{\prime}\oplus(\omega_{i,j}^{\prime})^{\vee}$
    given by the unramified $\operatorname{\mathcal{O}}_{R_{i}}$-action, and the splitting structure is reduced to a filtration
@@
    where each $\operatorname{\mathscr{F}}_{i,j}^{l}$ is locally a direct factor of rank $a_{i,j}^{1}+\cdots+a_{i,j}^{l}$, and the filtration satisfies the above condition (2).
@@
  • $\underline{\mathscr{F}_{\bullet}}=(\mathscr{F}_{i,j}^{l})$ is a splitting structure of $\underline{A}$.
  • The isomorphism is given by $\mathbb{Z}_{(p)}^{\times}$-isogenies between abelian schemes with $\operatorname{\mathcal{O}}_{B}$-structures and splitting structure.
@@
By Propositions 5.2 and 9.2 of [49], when $i$ is of type (AL) or type (C), such an isomorphism exists locally. Moreover, when $i$ is of type (AU), the same proof equally applies.
@@
which is $\prod_{i,j,l}G_{i,j}^{l}$-equivariant. The morphism $\pi$ is the natural forgetful morphism, which is a $\prod_{i,j,l}G_{i,j}^{l}$-torsor.
@@
is a resolution of singularities. Note that in this case, the associated parahoric subgroup $K_{p}=\operatorname{\mathcal{G}}(\mathbb{Z}_{p})$ is very special in the sense of [71] Definition 6.1. The purpose of this paper is to show that in our ramified setting, the smooth splitting model $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ admits many nice properties as in the unramified setting.
Recall that the Hodge cocharacter $\mu:\mathbb{G}_{m,F}\rightarrow G_{F}$ defines a parabolic subgroup $P_{\mu}=\{g\in G_{F}\,|\,\lim_{t\rightarrow 0}\mu(t)g\mu(t)^{-1}\,\textrm{exists}\}$ over $F$ in the usual way. For any $(V,\eta)\in\mathrm{Rep}_{F}P_{\mu}$, we have the associated automorphic vector bundle $\mathcal{V}$ over
@@
cf. [41]. Since $G$ is split over $F$, we have a canonical reductive model $\operatorname{\mathcal{O}}_{F}$, which by abuse of notation we still denote by $G$ (the precise meaning of the notation $G$ will be clear from the context; sometimes as in the last subsection we denote it by $\operatorname{\mathcal{G}}^{\mathrm{spl}}$). Moreover, the parabolic $P_{\mu}$ extends to a parabolic subgroup over $\operatorname{\mathcal{O}}_{F}$, which by abuse of notation we still denote by $P_{\mu}$, and $(V,\eta)$ naturally extends to a representation of $P_{\mu}$ over $\operatorname{\mathcal{O}}_{F}$. In general, it is hard to extend $\mathcal{V}$ to a vector bundle on $\operatorname{\mathscr{A}}_{\operatorname{\mathcal{O}}_{F}}$. Nevertheless, we have